# Quadratic Gating Functions in Mixture of Experts: A Statistical Insight

### v2

@@ Appendix D Proof of Theoretical Results
which indicates that $\|f_{G^{\prime}_{n}}-f_{G_{*}}\|_{L_{2}(\mu)}\to 0$ as $n\to\infty$. Recall that $\Theta$ is a compact set, therefore, we can replace the sequence $G^{\prime}_{n}$ by one of its subsequences that converges to a mixing measure $G^{\prime}\in\mathcal{G}_{N}(\Omega)$. Since $\mathcal{L}_{1}(G^{\prime}_{n},G_{*})>\varepsilon^{\prime}$, we deduce that $\mathcal{L}_{1}(G^{\prime},G_{*})>\varepsilon^{\prime}$.
@@
Thus, we get that $f_{G^{\prime}}({\bm{x}})=f_{G_{*}}({\bm{x}})$ for $\mu$-almost surely $x$. From Proposition E.1, we deduce that $G^{\prime}\equiv G_{*}$. Consequently, it follows that $\mathcal{L}_{1}(G^{\prime},G_{*})=0$, contradicting the fact that $\mathcal{L}_{1}(G^{\prime},G_{*})>\varepsilon^{\prime}>0$. Hence, the proof is completed.
D.3 Proof of Theorem C.1
In this proof, we first introduce the following lemma which will be used for our subsequent main proof of Theorem C.1.
Lemma D.2.
@@
$\lim_{\varepsilon\to 0}\inf_{G\in\mathcal{G}_{N}(\Theta):\mathcal{L}_{2,r}(G,G_{*})\leq\varepsilon}\frac{\|f_{G}-f_{G_{*}}\|_{L_{2}(\mu)}}{\mathcal{L}_{2,r}(G,G_{*})}=0.$ (54)
@@
$\inf_{\overline{G}_{n}\in\mathcal{G}_{N}(\Theta)}\sup_{G\in\mathcal{G}_{N}(\Theta)\setminus\mathcal{G}_{N^{*}-1}(\Theta)}\mathbb{E}_{f_{G}}[\mathcal{L}_{2,r}(\overline{G}_{n},G)]\gtrsim n^{-1/2},$ (55)
@@
Proof of Lemma D.2.
Firstly, note that from the Gaussian assumption on the noise variables, we obtain that $Y_{i}|X_{i}\sim\mathcal{N}(f_{G_{*}}({\bm{x}}_{i}),\sigma^{2})$ for all $i\in[n]$. Next, it follows from the assumption in equation (54) that for sufficiently small $\varepsilon>0$ and a fixed constant $C_{1}>0$ which we will choose later, there exists a mixing measure $G^{\prime}_{*}\in\mathcal{G}_{N}(\Theta)$ such that $\mathcal{L}_{2,r}(G^{\prime}_{*},G_{*})=2\varepsilon$ and $\|f_{G^{\prime}_{*}}-f_{G_{*}}\|_{L_{2}(\mu)}\leq C_{1}\varepsilon$. According to Le Cam’s lemma [43], since the Voronoi loss function $\mathcal{L}_{2,r}$ satisfies the weak triangle inequality, we get that
$\inf_{\overline{G}_{n}\in\mathcal{G}_{N}(\Theta)}\sup_{G\in\mathcal{G}_{N}(\Theta)\setminus\mathcal{G}_{N^{*}-1}(\Theta)}$ $\mathbb{E}_{f_{G}}[\mathcal{L}_{2,r}(\overline{G}_{n},G)]$
@@
Given the result of Lemma D.2, it suffices to prove that the following limit holds true for any $r\geq 1$:
$\lim_{\varepsilon\to 0}\inf_{G\in\mathcal{G}_{N}(\Theta):\mathcal{L}_{2,r}(G,G_{*})\leq\varepsilon}\frac{\|f_{G}-f_{G_{*}}\|_{L_{2}(\mu)}}{\mathcal{L}_{2,r}(G,G_{*})}=0.$ (57)
@@
As a result, $Q_{n}({\bm{x}})/\mathcal{L}_{2,r}(G_{n},G_{*})\to 0$ for $\mu$-almost surely $x$. Since the term $\sum_{j=1}^{N^{*}}\exp({\bm{x}}^{\top}{\bm{A}}_{j}^{*}{\bm{x}}+({\bm{b}}_{j}^{*})^{\top}{\bm{x}})$ is bounded, we deduce that $[f_{G_{n}}({\bm{x}})-f_{G_{*}}({\bm{x}})]/\mathcal{L}_{2,r}(G_{n},G_{*})\to 0$ for $\mu$-almost surely $x$. This result indicates that $\|f_{G_{n}}-f_{G_{*}}\|_{L_{2}(\mu)}/\mathcal{L}_{2,r}(G_{n},G_{*})\to 0$ as $n\to\infty$. Hence, we achieve the claim (57) and complete the proof.
D.4 Proof of Theorem B.1
The proof of Theorem B.1 can be done in a similar fashion to that of Theorem 3.1 in Appendix D.1.
D.5 Proof of Theorem B.3
@@
$\inf_{G\in\mathcal{G}_{N}(\Theta)}\|\tilde{f}_{G}-\tilde{f}_{G_{*}}\|_{L_{2}(\mu)}/\mathcal{L}_{3}(G,G_{*})>0.$ (59)
For that purpose, we divide the proof of the above inequality into local and global parts in the sequel. Here, we only present the proof of the local part, while that of the global part can be done using the same arguments as in Appendix D.2.
@@
$\lim_{\varepsilon\to 0}\inf_{G\in\mathcal{G}_{N}(\Theta):\mathcal{L}_{3}(G,G_{*})\leq\varepsilon}\|\tilde{f}_{G}-\tilde{f}_{G_{*}}\|_{L_{2}(\mu)}/\mathcal{L}_{3}(G,G_{*})>0.$ (60)
Assume by contrary that the above claim is not true true, then there exists a sequence of mixing measures $G_{n}=\sum_{i=1}^{N^{*}}\exp(c_{i}^{n})\delta_{({\bm{A}}^{n}_{i},{\bm{\eta}}^{n}_{i})}$ in $\mathcal{G}_{N}(\Theta)$ such that $\mathcal{L}_{3n}:=\mathcal{L}_{3}(G_{n},G_{*})\to 0$ and
@@
Step 1 - Taylor expansion. By abuse of notations, we sometimes tailor notations defined in Appendix D.2 to the setting of this proof. In this step, we would like to decompose the quantity
@@
into a combination of linearly independent elements using Taylor expansion. By using the same arguments for deriving equation (D.2), we get that $Q_{n}({\bm{x}})=A_{n}({\bm{x}})-B_{n}({\bm{x}})+C_{n}({\bm{x}})$, where
$A_{n}({\bm{x}})$ $:=\sum_{j=1}^{N^{*}}\sum_{i\in\mathcal{V}_{j}}\exp(c_{i}^{n})\Big{[}\exp({\bm{x}}^{\top}A_{i}^{n}{\bm{x}})h({\bm{x}};\eta_{i}^{n})-\exp({\bm{x}}^{\top}{\bm{A}}_{j}^{*}{\bm{x}})h({\bm{x}};{\bm{\eta}}_{j}^{*})\Big{]},$
@@
$C_{n}({\bm{x}})$ $:=\sum_{j=1}^{N^{*}}\Big{(}\sum_{i\in\mathcal{V}_{j}}\exp(c_{i}^{n})-\exp(c^{*}_{j})\Big{)}\Big{[}\exp({\bm{x}}^{\top}{\bm{A}}_{j}^{*}{\bm{x}})h({\bm{x}};{\bm{\eta}}_{j}^{*})-\exp({\bm{x}}^{\top}{\bm{A}}_{j}^{*}{\bm{x}})\tilde{f}_{G_{n}}({\bm{x}})\Big{]}.$
Decomposition of $A_{n}$. Let us denote $E({\bm{x}};A):=\exp({\bm{x}}^{\top}A{\bm{x}})$, then $A_{n}$ can be separated into two terms as follows:
$A_{n}({\bm{x}})$ $:=\sum_{j:|\mathcal{V}_{j}|=1}\sum_{i\in\mathcal{V}_{j}}\exp(c_{i}^{n})\Big{[}E({\bm{x}};A_{i}^{n})h({\bm{x}};\eta_{i}^{n})-E({\bm{x}};{\bm{A}}_{j}^{*})h({\bm{x}};{\bm{\eta}}_{j}^{*})\Big{]}$
$+\sum_{j:|\mathcal{V}_{j}|>1}\sum_{i\in\mathcal{V}_{j}}\exp(c_{i}^{n})\Big{[}E({\bm{x}};A_{i}^{n})h({\bm{x}};\eta_{i}^{n})-E({\bm{x}};{\bm{A}}_{j}^{*})h({\bm{x}};{\bm{\eta}}_{j}^{*})\Big{]}$
@@
$A_{n,1}({\bm{x}})$ $=\sum_{j:|\mathcal{V}_{j}|=1}\sum_{i\in\mathcal{V}_{j}}\frac{\exp(c_{i}^{n})}{\alpha!}\sum_{|\alpha|=1}(\Delta{\bm{A}}_{ij}^{n})^{\alpha_{1}}(\Delta{\bm{\eta}}_{ij}^{n})^{\alpha_{2}}\frac{\partial^{|\alpha_{1}|}E}{\partial A^{\alpha_{1}}}({\bm{x}};{\bm{A}}_{j}^{*})\frac{\partial^{|\alpha_{2}|}h}{\partial{\bm{\eta}}^{\alpha_{2}}}({\bm{x}};{\bm{\eta}}_{j}^{*})+R_{n,1}({\bm{x}})$
$=\sum_{j:|\mathcal{V}_{j}|=1}\sum_{|\alpha_{1}|+|\alpha_{2}|=1}S_{n,j,\alpha_{1},\alpha_{2}}\frac{\partial^{|\alpha_{1}|}E}{\partial A^{\alpha_{1}}}({\bm{x}};{\bm{A}}_{j}^{*})\frac{\partial^{|\alpha_{2}|}h}{\partial{\bm{\eta}}^{\alpha_{2}}}({\bm{x}};{\bm{\eta}}_{j}^{*})+R_{n,1}({\bm{x}}),$
@@
$A_{n,2}({\bm{x}})=\sum_{j:|\mathcal{V}_{j}|>1}\sum_{1\leq|\alpha_{1}|+|\alpha_{2}|\leq 2}S_{n,j,\alpha_{1},\alpha_{2}}\frac{\partial^{|\alpha_{1}|}E}{\partial A^{\alpha_{1}}}({\bm{x}};{\bm{A}}_{j}^{*})\frac{\partial^{|\alpha_{2}|}h}{\partial{\bm{\eta}}^{\alpha_{2}}}({\bm{x}};{\bm{\eta}}_{j}^{*})+R_{n,2}({\bm{x}}),$
@@
Decomposition of $B_{n}$. Recall that we have
@@
$Q_{n}({\bm{x}})$ $=\sum_{j=1}^{N^{*}}\sum_{0\leq|\alpha_{1}|+|\alpha_{2}|\leq 2}S_{n,j,\alpha_{1},\alpha_{2}}\frac{\partial^{|\alpha_{1}|}E}{\partial A^{\alpha_{1}}}({\bm{x}};{\bm{A}}_{j}^{*})\frac{\partial^{|\alpha_{2}|}h}{\partial{\bm{\eta}}^{\alpha_{2}}}({\bm{x}};{\bm{\eta}}_{j}^{*}),$
@@
with a note that at least one among them is non-zero. Then, from the decomposition of $Q_{n}({\bm{x}})$ in equation (D.5), we have
$\sum_{j=1}^{N^{*}}\sum_{|\alpha_{1}|+|\alpha_{2}|=0}^{1+\mathbf{1}_{\{|\mathcal{V}_{j}|>1\}}}\phi_{j,\alpha_{1},\alpha_{2}}\cdot$ $\frac{\partial^{|\alpha_{1}|}E}{\partial A^{\alpha_{1}}}({\bm{x}};{\bm{A}}_{j}^{*})\frac{\partial^{|\alpha_{2}|}h}{\partial{\bm{\eta}}^{\alpha_{2}}}({\bm{x}};{\bm{\eta}}_{j}^{*}),$
@@
for $\mu$-almost surely $x$. It is worth noting that the term $\frac{\partial^{|\alpha_{1}|}E}{\partial A^{\alpha_{1}}}({\bm{x}};{\bm{A}}_{j}^{*})\frac{\partial^{|\alpha_{2}|}h}{\partial{\bm{\eta}}^{\alpha_{2}}}({\bm{x}};{\bm{\eta}}_{j}^{*})$ can be explicitly expressed as
  • When $|\alpha_{1}|=0,|\alpha_{2}|=0$: $\exp({\bm{x}}^{\top}{\bm{A}}_{j}^{*}{\bm{x}})h({\bm{x}};{\bm{\eta}}_{j}^{*})$;
  • When $|\alpha_{1}|=1,|\alpha_{2}|=0$: $x^{(u)}x^{(v)}\exp({\bm{x}}^{\top}{\bm{A}}_{j}^{*}{\bm{x}})h({\bm{x}};{\bm{\eta}}_{j}^{*})$;
@@
  • When $|\alpha_{1}|=2,|\alpha_{2}|=0$: $x^{(u)}x^{(v)}x^{(u^{\prime})}x^{(v^{\prime})}\exp({\bm{x}}^{\top}{\bm{A}}_{j}^{*}{\bm{x}})h({\bm{x}};{\bm{\eta}}_{j}^{*})$;
@@
Recall that the expert function $h$ satisfies the condition in Definition B.2, i.e. the set
@@
Appendix E Additional Results
In this appendix, we study the identifiability of the MoE models with the quadratic polynomial gate and the quadratic monomial gate in that order.
Proposition E.1.

### v3

@@ Appendix A Proof of Theoretical Results
which indicates that $\|f_{G^{\prime}_{n}}-f_{G_{*}}\|_{L_{2}(\mu)}\to 0$ as $n\to\infty$. Recall that $\Theta$ is a compact set, therefore, we can replace the sequence $G^{\prime}_{n}$ by one of its subsequences that converges to a mixing measure $G^{\prime}\in\mathcal{M}_{N}(\Omega)$. Since $\mathcal{L}_{1}(G^{\prime}_{n},G_{*})>\varepsilon^{\prime}$, we deduce that $\mathcal{L}_{1}(G^{\prime},G_{*})>\varepsilon^{\prime}$.
@@
Thus, we get that $f_{G^{\prime}}({\bm{x}})=f_{G_{*}}({\bm{x}})$ for $\mu$-almost surely $x$. From Proposition B.1, we deduce that $G^{\prime}\equiv G_{*}$. Consequently, it follows that $\mathcal{L}_{1}(G^{\prime},G_{*})=0$, contradicting the fact that $\mathcal{L}_{1}(G^{\prime},G_{*})>\varepsilon^{\prime}>0$. Hence, the proof is completed.
A.3 Proof of Theorem 3.4
In this proof, we first introduce the following lemma which will be used for our subsequent main proof of Theorem 3.4.
Lemma A.2.
@@
$\lim_{\varepsilon\to 0}\inf_{G\in\mathcal{M}_{N}(\Theta):\mathcal{L}_{2,r}(G,G_{*})\leq\varepsilon}\frac{\|f_{G}-f_{G_{*}}\|_{L_{2}(\mu)}}{\mathcal{L}_{2,r}(G,G_{*})}=0.$ (53)
@@
$\inf_{\overline{G}_{n}\in\mathcal{M}_{N}(\Theta)}\sup_{G\in\mathcal{M}_{N}(\Theta)\setminus\mathcal{M}_{N^{*}-1}(\Theta)}\mathbb{E}_{f_{G}}[\mathcal{L}_{2,r}(\overline{G}_{n},G)]\gtrsim n^{-1/2},$ (54)
@@
Proof of Lemma A.2.
Firstly, note that from the Gaussian assumption on the noise variables, we obtain that $Y_{i}|X_{i}\sim\mathcal{N}(f_{G_{*}}({\bm{x}}_{i}),\sigma^{2})$ for all $i\in[n]$. Next, it follows from the assumption in equation (53) that for sufficiently small $\varepsilon>0$ and a fixed constant $C_{1}>0$ which we will choose later, there exists a mixing measure $G^{\prime}_{*}\in\mathcal{M}_{N}(\Theta)$ such that $\mathcal{L}_{2,r}(G^{\prime}_{*},G_{*})=2\varepsilon$ and $\|f_{G^{\prime}_{*}}-f_{G_{*}}\|_{L_{2}(\mu)}\leq C_{1}\varepsilon$. According to Le Cam’s lemma yu97lecam , since the Voronoi loss function $\mathcal{L}_{2,r}$ satisfies the weak triangle inequality, we get that
$\inf_{\overline{G}_{n}\in\mathcal{M}_{N}(\Theta)}$ $\sup_{G\in\mathcal{M}_{N}(\Theta)\setminus\mathcal{M}_{N^{*}-1}(\Theta)}\mathbb{E}_{f_{G}}[\mathcal{L}_{2,r}(\overline{G}_{n},G)]$
@@
Given the result of Lemma A.2, it suffices to prove that the following limit holds true for any $r\geq 1$:
$\lim_{\varepsilon\to 0}\inf_{G\in\mathcal{M}_{N}(\Theta):\mathcal{L}_{2,r}(G,G_{*})\leq\varepsilon}\frac{\|f_{G}-f_{G_{*}}\|_{L_{2}(\mu)}}{\mathcal{L}_{2,r}(G,G_{*})}=0.$ (56)
@@
As a result, $Q_{n}({\bm{x}})/\mathcal{L}_{2,r}(G_{n},G_{*})\to 0$ as $n\to\infty$ for $\mu$-almost surely $x$. Since the term $\sum_{j=1}^{N^{*}}\exp({\bm{x}}^{\top}{\bm{A}}_{j}^{*}{\bm{x}}+({\bm{b}}_{j}^{*})^{\top}{\bm{x}})$ is bounded, we deduce that $[f_{G_{n}}({\bm{x}})-f_{G_{*}}({\bm{x}})]/\mathcal{L}_{2,r}(G_{n},G_{*})\to 0$ for $\mu$-almost surely $x$. This result indicates that $\|f_{G_{n}}-f_{G_{*}}\|_{L_{2}(\mu)}/\mathcal{L}_{2,r}(G_{n},G_{*})\to 0$ as $n\to\infty$. Hence, we achieve the claim (56) and complete the proof.
A.4 Proof of Theorem 3.5
The proof of Theorem 3.5 can be done in a similar fashion to that of Theorem 3.1 in Appendix A.1.
A.5 Proof of Theorem 3.7
@@
$\inf_{G\in\mathcal{M}_{N}(\Theta)}\|\tilde{f}_{G}-\tilde{f}_{G_{*}}\|_{L_{2}(\mu)}/\mathcal{L}_{3}(G,G_{*})>0.$ (58)
For that purpose, we divide the proof of the above inequality into local and global parts in the sequel. Here, we only present the proof of the local part, while that of the global part can be done using the same arguments as in Appendix A.2.
@@
$\lim_{\varepsilon\to 0}\inf_{G\in\mathcal{M}_{N}(\Theta):\mathcal{L}_{3}(G,G_{*})\leq\varepsilon}\|\tilde{f}_{G}-\tilde{f}_{G_{*}}\|_{L_{2}(\mu)}/\mathcal{L}_{3}(G,G_{*})>0.$ (59)
Assume by contrary that the above claim is not true true, then there exists a sequence of mixing measures $G_{n}=\sum_{i=1}^{N^{*}}\exp(c_{i}^{n})\delta_{({\bm{A}}^{n}_{i},{\bm{\eta}}^{n}_{i})}$ in $\mathcal{M}_{N}(\Theta)$ such that $\mathcal{L}_{3n}:=\mathcal{L}_{3}(G_{n},G_{*})\to 0$ and
@@
Step 1 - Taylor expansion. By abuse of notations, we sometimes tailor notations defined in Appendix A.2 to the setting of this proof. In this step, we would like to decompose the quantity
@@
into a combination of linearly independent elements using Taylor expansion. By using the same arguments for deriving equation (A.2), we get that $Q_{n}({\bm{x}})=A_{n}({\bm{x}})-B_{n}({\bm{x}})+C_{n}({\bm{x}})$, where
$A_{n}({\bm{x}})$ $:=\sum_{j=1}^{N^{*}}\sum_{i\in\mathcal{V}_{j}}\exp(c_{i}^{n})\Big{[}\exp({\bm{x}}^{\top}A_{i}^{n}{\bm{x}})\mathcal{E}({\bm{x}};\eta_{i}^{n})-\exp({\bm{x}}^{\top}{\bm{A}}_{j}^{*}{\bm{x}})\mathcal{E}({\bm{x}};{\bm{\eta}}_{j}^{*})\Big{]},$
@@
$C_{n}({\bm{x}})$ $:=\sum_{j=1}^{N^{*}}\Big{(}\sum_{i\in\mathcal{V}_{j}}\exp(c_{i}^{n})-\exp(c^{*}_{j})\Big{)}\Big{[}\exp({\bm{x}}^{\top}{\bm{A}}_{j}^{*}{\bm{x}})\mathcal{E}({\bm{x}};{\bm{\eta}}_{j}^{*})-\exp({\bm{x}}^{\top}{\bm{A}}_{j}^{*}{\bm{x}})\tilde{f}_{G_{n}}({\bm{x}})\Big{]}.$
Decomposition of $A_{n}({\bm{x}})$. Let us denote $E({\bm{x}};A):=\exp({\bm{x}}^{\top}A{\bm{x}})$, then $A_{n}({\bm{x}})$ can be separated into two terms as follows:
$A_{n}({\bm{x}})$ $:=\sum_{j:|\mathcal{V}_{j}|=1}\sum_{i\in\mathcal{V}_{j}}\exp(c_{i}^{n})\Big{[}E({\bm{x}};A_{i}^{n})\mathcal{E}({\bm{x}};\eta_{i}^{n})-E({\bm{x}};{\bm{A}}_{j}^{*})\mathcal{E}({\bm{x}};{\bm{\eta}}_{j}^{*})\Big{]}$
$+\sum_{j:|\mathcal{V}_{j}|>1}\sum_{i\in\mathcal{V}_{j}}\exp(c_{i}^{n})\Big{[}E({\bm{x}};A_{i}^{n})\mathcal{E}({\bm{x}};\eta_{i}^{n})-E({\bm{x}};{\bm{A}}_{j}^{*})\mathcal{E}({\bm{x}};{\bm{\eta}}_{j}^{*})\Big{]}$
@@
$A_{n,1}({\bm{x}})$ $=\sum_{j:|\mathcal{V}_{j}|=1}\sum_{i\in\mathcal{V}_{j}}\frac{\exp(c_{i}^{n})}{\alpha!}\sum_{|\alpha|=1}(\Delta{\bm{A}}_{ij}^{n})^{\alpha_{1}}(\Delta{\bm{\eta}}_{ij}^{n})^{\alpha_{2}}\frac{\partial^{|\alpha_{1}|}E}{\partial A^{\alpha_{1}}}({\bm{x}};{\bm{A}}_{j}^{*})\frac{\partial^{|\alpha_{2}|}\mathcal{E}}{\partial{\bm{\eta}}^{\alpha_{2}}}({\bm{x}};{\bm{\eta}}_{j}^{*})+R_{n,1}({\bm{x}})$
$=\sum_{j:|\mathcal{V}_{j}|=1}\sum_{|\alpha_{1}|+|\alpha_{2}|=1}S_{n,j,\alpha_{1},\alpha_{2}}\frac{\partial^{|\alpha_{1}|}E}{\partial A^{\alpha_{1}}}({\bm{x}};{\bm{A}}_{j}^{*})\frac{\partial^{|\alpha_{2}|}\mathcal{E}}{\partial{\bm{\eta}}^{\alpha_{2}}}({\bm{x}};{\bm{\eta}}_{j}^{*})+R_{n,1}({\bm{x}}),$
@@
$A_{n,2}({\bm{x}})=\sum_{j:|\mathcal{V}_{j}|>1}\sum_{1\leq|\alpha_{1}|+|\alpha_{2}|\leq 2}S_{n,j,\alpha_{1},\alpha_{2}}\frac{\partial^{|\alpha_{1}|}E}{\partial A^{\alpha_{1}}}({\bm{x}};{\bm{A}}_{j}^{*})\frac{\partial^{|\alpha_{2}|}\mathcal{E}}{\partial{\bm{\eta}}^{\alpha_{2}}}({\bm{x}};{\bm{\eta}}_{j}^{*})+R_{n,2}({\bm{x}}),$
@@
Decomposition of $B_{n}({\bm{x}})$. Recall that we have
@@
$Q_{n}({\bm{x}})$ $=\sum_{j=1}^{N^{*}}\sum_{0\leq|\alpha_{1}|+|\alpha_{2}|\leq 2}S_{n,j,\alpha_{1},\alpha_{2}}\frac{\partial^{|\alpha_{1}|}E}{\partial A^{\alpha_{1}}}({\bm{x}};{\bm{A}}_{j}^{*})\frac{\partial^{|\alpha_{2}|}\mathcal{E}}{\partial{\bm{\eta}}^{\alpha_{2}}}({\bm{x}};{\bm{\eta}}_{j}^{*}),$
@@
with a note that at least one among them is non-zero. Then, from the decomposition of $Q_{n}({\bm{x}})$ in equation (A.5), we have
$\sum_{j=1}^{N^{*}}\sum_{|\alpha_{1}|+|\alpha_{2}|=0}^{1+\mathbf{1}_{\{|\mathcal{V}_{j}|>1\}}}\phi_{j,\alpha_{1},\alpha_{2}}\cdot$ $\frac{\partial^{|\alpha_{1}|}E}{\partial A^{\alpha_{1}}}({\bm{x}};{\bm{A}}_{j}^{*})\frac{\partial^{|\alpha_{2}|}\mathcal{E}}{\partial{\bm{\eta}}^{\alpha_{2}}}({\bm{x}};{\bm{\eta}}_{j}^{*}),$
@@
for $\mu$-almost surely $x$. It is worth noting that the term $\frac{\partial^{|\alpha_{1}|}E}{\partial A^{\alpha_{1}}}({\bm{x}};{\bm{A}}_{j}^{*})\frac{\partial^{|\alpha_{2}|}\mathcal{E}}{\partial{\bm{\eta}}^{\alpha_{2}}}({\bm{x}};{\bm{\eta}}_{j}^{*})$ can be explicitly expressed as
  • When $|\alpha_{1}|=0,|\alpha_{2}|=0$: $\exp({\bm{x}}^{\top}{\bm{A}}_{j}^{*}{\bm{x}})\mathcal{E}({\bm{x}};{\bm{\eta}}_{j}^{*})$;
  • When $|\alpha_{1}|=1,|\alpha_{2}|=0$: $x^{(u)}x^{(v)}\exp({\bm{x}}^{\top}{\bm{A}}_{j}^{*}{\bm{x}})\mathcal{E}({\bm{x}};{\bm{\eta}}_{j}^{*})$;
@@
  • When $|\alpha_{1}|=2,|\alpha_{2}|=0$: $x^{(u)}x^{(v)}x^{(u^{\prime})}x^{(v^{\prime})}\exp({\bm{x}}^{\top}{\bm{A}}_{j}^{*}{\bm{x}})\mathcal{E}({\bm{x}};{\bm{\eta}}_{j}^{*})$;
@@
Recall that the expert function $h$ satisfies the condition in Definition 3.6, i.e. the set
@@
Appendix B Identifiability of Quadratic Gating MoE
In this appendix, we study the identifiability of the MoE models with the quadratic polynomial gate and the quadratic monomial gate in Proposition B.1 and Proposition B.2, respectively.
Proposition B.1.